OpenStudy (anonymous):

factorise x-7x-10 =0 thanks :)

OpenStudy (anonymous):

is it x^2-7x-10=0 ?

OpenStudy (anonymous):

yes it is x^2

OpenStudy (anonymous):

It would have been easy if it was x^2-7x+10=0.

OpenStudy (anonymous):

well it is a -10!!!!

OpenStudy (anonymous):

well, then we can solve the equation using x = \[-b\pm\sqrt{b ^{2}-4 \times a \times c}\]\[\div (2 \times a)\] we get x = \[(7 \pm \sqrt{89})/2\]

OpenStudy (anonymous):

oh thanks alot im still working on it